OpenStudy (anonymous):

g(x)=[(ax+b)^(1/2)-1]/x for x not equal to 0 and g(x)=1 for x=0 if g(x) is continuous at x=0 then a+b=?

OpenStudy (amistre64):

Lhop perhaps?

OpenStudy (anonymous):

but the answer is given 3 i got 2

OpenStudy (amistre64):

\[\lim\frac{(ax+b)^{1/2}-1}{x}\] \[\lim\frac{a}{2(ax+b)^{1/2}}\] \[\lim\frac{a}{2b^{1/2}}=1\] a = 2sqrt(b); or b = a^2/4 did i mess that up any?

OpenStudy (anonymous):

well what after that

OpenStudy (amistre64):

\[\frac{(2b^{1/2}x+b)^{1/2}-1}{x}\] hmm, or maybe \[\frac{(ax+b)^{1/2}-1}{x}~if~(ax+b)^{1/2}=x+1\]

OpenStudy (amistre64):

\[ax+b = x^2+2x+1\] \[b = x^2+(2-a)x+1\] \[b+a = x^2+(2-a)x+1+a\] at x=0, b+a = 1+a; b=1 if b=1, then from the first run, b = a^2/4, if |a| = 2

OpenStudy (amistre64):

so we could test out: b = 1, a = -2, or 2

OpenStudy (anonymous):

thanks a lot

OpenStudy (amistre64):

youre welcome ... how did you try it?

OpenStudy (anonymous):

well i was stuck after applying l'hos

OpenStudy (amistre64):

so was i :) i spose since a=2sqrt(b), and b=1 .. a=2 i just had a hunch to try to make the top equal to x is all